# Residual permutation test for regression coefficient testing

### v3

@@ Abstract
We consider the problem of testing whether a single coefficient is equal to zero in linear models when the dimension of covariates $p$ can be up to a constant fraction of sample size $n$. In this regime, an important topic is to propose tests with finite-population valid size control without requiring the noise to follow strong distributional assumptions. In this paper, we propose a new method, called residual permutation test (RPT), which is constructed by projecting the regression residuals onto the space orthogonal to the union of the column spaces of the original and permuted design matrices. RPT can be proved to achieve finite-population size validity under fixed design with just exchangeable noises, whenever $p<n/2$. Moreover, RPT is shown to be asymptotically powerful for heavy tailed noises with bounded $(1+t)$-th order moment when the true coefficient is at least of order $n^{-t/(1+t)}$ for $t\in[0,1]$. We further proved that this signal size requirement is essentially rate-optimal in the minimax sense. Numerical studies confirm that RPT performs well in a wide range of simulation settings with normal and heavy-tailed noise distributions.
Keywords: permutation test, finite-population validity, heavy tail distribution, high-dimensional data
@@ 1 Introduction
Here, we are primarily interested in designing a new coefficient test with finite-population size validity. In other words, we require our test to have valid size control with arbitrary magnitude of $n$, instead of requiring some asymptotic regime assumption that may be unrealistic in practice. Classically, one of the most well-known tests for $b=0$ in model (1) is the ANOVA test (Fisher, 1935). It is known to be asymptotically valid when $p$ is fixed and $n\to\infty$, and has finite-population valid size control when the noise possesses spherical symmetry. However, as we will see in Section 3, the finite-population size of ANOVA test can be far from the nominal level in the presence of heavy-tailed noises. This motivates us to propose a new test that has finite-population valid size control under weaker noise assumptions, especially with relatively large $p$, which is also a topic that has received increasing attention in recent years (see e.g. Section 2). In particular, we are interested in developing a test with finite-population valid size control under a fixed design of $\boldsymbol{X}$ and $\boldsymbol{Z}$ by assuming that the noise $\boldsymbol{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{n})^{\top}$ has exchangeable components:
@@
We remark that Assumption 1 differs from the common assumption that $\mathbb{E}[\boldsymbol{\varepsilon}]=\boldsymbol{0}$ for fixed design of $\boldsymbol{X},\boldsymbol{Z}$, or that $\mathbb{E}[\boldsymbol{\varepsilon}\mid\boldsymbol{X},\boldsymbol{Z}]=0$ (or even the more relaxed assumption $\mathbb{E}[\boldsymbol{\varepsilon}^{\top}(\boldsymbol{X},\boldsymbol{Z})]=0$) when $\boldsymbol{X},\boldsymbol{Z}$ follow a random design, which allows for heteroskedastic noise and typically appears in the regression coefficient tests with asymptotically valid size control. A common approach to handle exchangeable noise is through the idea of permutation tests (Pitman, 1937a, b, 1938). Recently, Lei and Bickel (2021) implemented this idea to the problem of regression coefficient testing. In their seminal work, the authors proposed a cyclic permutation test that achieved finite population size validity under Assumption 1 by exploiting the exchangeability of the noise terms. However, to achieve a size $\alpha$ control, their cyclic permutation test requires that $n/p\geq 1/\alpha-1$. For instance, for a sample size of $n=300$ and a targeting Type-I error rate is $\alpha=0.01$, at most $p=2$ covariates are allowed in $\boldsymbol{X}$. This limits the applicability of their test in large dimensions. In this paper, we consider the more challenging question of finite-population Type-I error control in setting where $p$ is allowed to be of the same order of magnitude as $n$. We propose a residual permutation test (RPT), a permutation-based approach that performs hypothesis tests by manipulating the empirical residuals after regression adjustment. The proposed test is guaranteed to have the correct Type-I error control whenever $p<n/2$. Moreover, our result is fixed design and does not require any regularity conditions on the design matrix $\boldsymbol{X}$.
@@
In addition to proving its finite-population size validity, we further analyze the statistical power of the proposed test in the regime where $p$ can be up to a constant fraction of $n$, which we will refer to as the proportional regime in this work, especially when the $\varepsilon_{i}$’s follow a heavy-tailed distribution. It should be noted that just with Assumption 1 we can only guarantee our test to have correct size control, the resulting test may not necessarily have power. Indeed, just with Assumption 1, $b$ may not even be identifiable, so that there is no test that is uniformly valid under the null while still maintainng power against some alternative. For example, in the extreme case where $\boldsymbol{Z}=\boldsymbol{X}\beta^{Z}$ for some $\beta^{Z}\in\mathbb{R}^{p}$, i.e., that $\boldsymbol{Z}$ can be expressed as a linear representation of the column vectors of $\boldsymbol{X}$, then $b$ is not an identifiable parameter anymore, i.e., no test can have any power. As we can also see in (6), under Assumption 1 and with such choice of $\boldsymbol{Z}$, we can always have $\mathbb{P}(\phi\leq\alpha)\leq\alpha$ for all $\alpha\in[0,1]$, no matter how large $b$ is. This means that in order for our test to have power, additional assumption on $\boldsymbol{Z}$ must be imposed. As we will discuss further in Section 2.3, statistical methods with robustness to heavy-tailed data have significant demands in practice (Eklund, Nichols and Knutsson, 2016; Wang, Peng and Li, 2015; Cont, 2001), and has been actively studied in both modern statistics and theoretical computer science communities. Despite its importance, there is a lack of available tools that can handle regression coefficient testing under this proportional regime with heavy-tailed noise. In this paper, we fill this gap by showing that under a suitable modelling assumption of $\boldsymbol{Z}$, when the $\varepsilon_{i}$’s are i.i.d. and have a finite $(1+t)$-th order moment for any $t\in[0,1]$, and that $n/p\geq 3+m$ for some $m>0$, our proposed test is asymptotically powerful whenever the coefficient $b$ is of order at least $n^{-t/(1+t)}$. In proving this result, a crucial step is to establish a concentration bound for projected length of a random vector with independent heavy-tailed components. This concentration bound may be of independent interest for future research on statistical procedures with heavy-tailed noise, and is stated in Corollary 8. We also studied the minimax rate optimality of regression coefficient testing with heavy-tailed noises; and proved that in the presence of heavy-tailed noise with only a finite $(1+t)$-th moment, the $n^{-t/(1+t)}$ order requirement for $b$ is essentially rate-optimal.
@@
Since ANOVA has been used extensively in practical applications, as an independent contribution, we provide a more comprehensive analysis of the ANOVA test. Specifically, while ANOVA can be shown to have finite-population size validity with spherically symmetric noise, our simulations show that the it can substantially violate the nominal size control under more general noise distributions. At the same, we propose another permutation-based test: naive residual permutation test (naive RPT), which like ANOVA, also has valid size control under spherically symmetric noise distribution whenever $p<n$. While naive RPT is still not valid for non-spherically symmetric noises, it does appear to have smaller Type I error violations compared to ANOVA.
@@
  • We propose a new test that has finite population size validity with fixed-design linear models and exchangeable noises whenever $p<n/2$.
@@
The rest of this paper is organized as follows. In Section 2, we review existing results in regression coefficient testing, permutation- and randomization-based tests and heavy-tailed data. In Section 3, we provide more studies on the finite-sample properties of ANOVA test with non-Gaussian noises, and propose a new test that is easier to implement and more robust to non-Gaussianity. As ANOVA test has been heavily used in practical applications, we believe this is of independent interest. In Section 4, we present our method, and prove its finite population size validity. In Sections 5 and 7, we provide power analysis of RPT and study its minimax rate optimality under some heavy-tailed assumptions. Finally, in Section 8 we provide numerical analysis. In Section 9, we end the manuscript with a discussion.
@@ 2.1 Hypothesis testing of regression coefficients
As emphasized by Lei and Bickel (2021), this is a challenging problem, with a “century long effort” in the statistical community to alleviate the strong assumption of ANOVA. In the context of finite-population size validity, some representative works include Hartigan (1970); Meinshausen (2015). However, the two methods mentioned above still require the noise to follow certain geometric constraint, which is either symmetric about 0 or rotationally invariant. Lei and Bickel (2021) represented, to the best of our knowledge, the first work that established finite-population size control with only exchangeable noise. However, as mentioned in the introduction, the cyclic permutation test proposed in Lei and Bickel (2021) requires the dimension of $p$ to be much smaller than $n$ for valid size control, and no corresponding statistical power analysis was provided. Alternative test with less restrictive assumptions on dimension $p$ was proposed in D’Haultfœuille and Tuvaandorj (2022), which relaxes the dimensionality assumption by requiring the rows of $\boldsymbol{X}$ to follow a discrete random distribution with a relatively small number of unique values.
Besides finite-population size validity, a less demanding criteria for coefficient test is the asymptotic size validity. The idea of permutation or randomization have been heavily used to propose asymptotically valid test; see Section 2.2 for more details. In the high-dimensional regime where $p$ is proportional or even much larger than $n$, debiased / desparsified Lasso was proposed to construct confidence intervals and perform coefficient tests (Zhang and Zhang, 2014; Van de Geer et al., 2014; Javanmard and Montanari, 2014). By invoking 1) certain sparsity conditions on the regression coefficients; 2) some regularity conditions on the design matrix $\boldsymbol{X}$ and 3) sharp tail bounds on the noise variables, debiased / desparsified Lasso is guaranteed to establish asymptotically valid p-value and confidence intervals for regression coefficients. We remark that the additional sparsity assumption on the regression coefficients allow for the dimension $p$ to diverge at a much faster rate than $n$ compared to asymptotic regime studied in the current paper. Other follow up studies include Zhu and Bradic (2018); Bradic et al. (2019); Shah and Bühlmann (2019), to name a few.
@@ 2.2 Permutation- and randomization-based hypothesis tests
Other related applications of permutation tests include sharp null hypothesis tests (Caughey, Dafoe and Miratrix, 2017; Caughey et al., 2021), instrumental variable tests (Imbens and Rosenbaum, 2005), and conditional independence tests (Berrett et al., 2020; Kim et al., 2021).
@@ 2.3 Heavy-tailed data
Since the pioneering work by Catoni (2012), the problem of extracting useful information from heavy-tailed data (or the related adversarially contaminated data) has been an active area of research in mathematical statistics and theoretical computer science literature in the past ten years (Bubeck, Cesa-Bianchi and Lugosi, 2013; Lykouris, Mirrokni and Paes Leme, 2018; Lugosi and Mendelson, 2019; Sun, Zhou and Fan, 2020; Fan, Wang and Zhu, 2021). When we allow the dimension $p$ to grow with $n$, heavy-tailed data has been actively studied in mean estimation  (Lugosi and Mendelson, 2019, 2021), regression coefficient estimation (Wang, 2013; Fan, Li and Wang, 2017; Sun, Zhou and Fan, 2020; Pensia, Jog and Loh, 2020) and covariance matrix analysis (Loh and Tan, 2018; Fan, Wang and Zhu, 2021). The definition of “heavy-tail” may vary across different articles. Among all literature working with heavy-tailed noise, our assumptions are most similar to those in Sun, Zhou and Fan (2020); Bubeck, Cesa-Bianchi and Lugosi (2013), which assume that the noise variables has at most a finite $(1+t)$-th order moments for some $t\in(0,1]$ without any geometric or shape constraints. To our knowledge this is also the weakest heavy tail assumption studied in the literature.
@@
3 Finite-population size validity of ANOVA beyond Gaussianity
@@
Table 1 summarizes the performance of ANOVA test from $100000$ Monte Carlo simulations. We consider the sizes of the ANOVA test at nominal levels $\alpha=0.01,0.05$. According to the simulation results, when the noises of $\boldsymbol{e}$ and $\boldsymbol{\varepsilon}$ follows a standard normal distribution, the ANOVA test has the correct size control, which is consistent with Lemma 1. However, when normality is violated, the ANOVA test will be overly optimistic, with an empirical size more than twice as large as the nominal level in some $1\%$-level tests (this issue is more pronounced if we consider a $0.5\%$ test level; see Table A2 in the appendix). In particular, the performance of noise type $t_{1}$ is in general worse than that of $t_{2}$, this means that ANOVA test is more vulnerable to heavy-tailed noises. Moreover, the performance of ANOVA is worse with a heavy-tailed design matrix $\boldsymbol{X}$.
@@
It is worth noting that when $\beta=0$ in (1), we can easily construct a permutation test with valid size control by comparing the correlation of $Y$ to $Z$ and to its permutations. From this intuition, a straightforward approach is to first regress both $\boldsymbol{Y}$ and $\boldsymbol{Z}$ onto $\boldsymbol{X}$ to eliminate the influence of $\boldsymbol{X}$, and then to use regression residuals for permutation test construction. Specifically, let $\hat{\boldsymbol{R}}_{\varepsilon}:=(\boldsymbol{I}-\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top})\boldsymbol{Y}$ and $\hat{\boldsymbol{R}}_{e}:=(\boldsymbol{I}-\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top})\boldsymbol{Z}$ be the regression residuals after projecting $\boldsymbol{Y}$ and $\boldsymbol{Z}$ onto $\boldsymbol{X}$ respectively. Let $\boldsymbol{V}_{0}\in\mathbb{R}^{n\times(n-p)}$ be a matrix with orthonormal columns spanning an $(n-p)$-dimensional subspace of $\mathrm{span}(\boldsymbol{X})^{\perp}$, then $\boldsymbol{I}-\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}=\boldsymbol{V}_{0}\boldsymbol{V}_{0}^{\top}$. Hence under $H_{0}:b=0$, the regression residuals $\hat{\boldsymbol{R}}_{\varepsilon}$ satisfy $\hat{\boldsymbol{R}}_{\varepsilon}=\boldsymbol{V}_{0}\boldsymbol{V}_{0}^{\top}\boldsymbol{Y}=\boldsymbol{V}_{0}\boldsymbol{V}_{0}^{\top}\boldsymbol{\varepsilon}$. From above, we construct a test, which we call as naive residual permutation test, based on the projected residuals $\hat{\boldsymbol{\varepsilon}}:=\boldsymbol{V}_{0}^{\top}\hat{\boldsymbol{R}}_{\varepsilon}=\boldsymbol{V}_{0}^{\top}\boldsymbol{Y}$ and $\hat{e}:=\boldsymbol{V}_{0}^{\top}\hat{\boldsymbol{R}}_{e}=\boldsymbol{V}_{0}^{\top}\boldsymbol{Z}$ as
@@
where the $\boldsymbol{P}_{k}\in\mathbb{R}^{(n-p)\times(n-p)}$’s are random permutation matrices that are sampled uniformly at random from the set of all permutation matrices. Lemma 2 shows that under a slightly weaker condition than Lemma 1, $\phi_{\mathrm{naive}}$ has valid Type I error control.
@@
While Lemma 2 is slightly less stringent than Lemma 1, it still requires the spherical symmetry in distributions. To better understand their empirical performances, we also show the performance of $\phi_{\mathrm{naive}}$ with non-Gaussian noises or non-Gaussian designs in Table 1 and Figures 1(d)-(f) . Without the strong Gaussianity or spherically symmetry assumption, $\phi_{\mathrm{anova}}$ is also not guaranteed to have finite-population size validity. Nevertheless, when both tests are invalid, the size of naive permutation test is closer to the correct level than its competitor. This indicates that naive test is more robust to non-Gaussian distributions. Moreover, the naive test is an intuitive method and is easy to implement. Thus, the naive test could be used as a preferrable alternative to ANOVA in real data analysis when $n/2\leq p<n$.

### v4

@@ Abstract
We consider the problem of testing whether a single coefficient is equal to zero in linear models when the dimension of covariates $p$ can be up to a constant fraction of sample size $n$. In this regime, an important topic is to propose tests with finite-sample valid size control without requiring the noise to follow strong distributional assumptions. In this paper, we propose a new method, called residual permutation test (RPT), which is constructed by projecting the regression residuals onto the space orthogonal to the union of the column spaces of the original and permuted design matrices. RPT can be proved to achieve finite-sample size validity under fixed design with just exchangeable noises, whenever $p<n/2$. Moreover, RPT is shown to be asymptotically powerful for heavy tailed noises with bounded $(1+t)$-th order moment when the true coefficient is at least of order $n^{-t/(1+t)}$ for $t\in[0,1]$. We further proved that this signal size requirement is essentially rate-optimal in the minimax sense. Numerical studies confirm that RPT performs well in a wide range of simulation settings with normal and heavy-tailed noise distributions.
62J99,
62F03,
permutation test,
finite-sample validity,
heavy tail distribution,
high-dimensional data,
keywords:
[class=MSC]
keywords:
[…]
thanks:  Equal contributionthanks:  Part of this work was done while KW was at the Institute for Interdisciplinary Information Sciences, Tsinghua University.
, 11footnotemark: 1 and thanks: To whom correspondence should be addressed
@@ 1 Introduction
Here, we are primarily interested in designing a new coefficient test with finite-sample size validity. In other words, we require our test to have valid size control with arbitrary magnitude of $n$, instead of requiring some asymptotic regime assumption that may be unrealistic in practice. Classically, one of the most well-known tests for $b=0$ in model (1) is the ANOVA test (Fisher, 1935). It is known to be asymptotically valid when $p$ is fixed and $n\to\infty$, and has finite-sample valid size control when the noise possesses spherical symmetry. However, as we will see in Section 3, the finite-sample size of ANOVA test can be far from the nominal level in the presence of heavy-tailed noises. This motivates us to propose a new test that has finite-sample valid size control under weaker noise assumptions, especially with relatively large $p$, which is also a topic that has received increasing attention in recent years (see e.g. Section 2). In particular, we are interested in developing a test with finite-sample valid size control under a fixed design of $\boldsymbol{X}$ and $\boldsymbol{Z}$ by assuming that the noise $\boldsymbol{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{n})^{\top}$ has exchangeable components:
@@
We remark that Assumption 1 differs from the common assumption that $\mathbb{E}[\boldsymbol{\varepsilon}]=\boldsymbol{0}$ for fixed design of $\boldsymbol{X},\boldsymbol{Z}$, or that $\mathbb{E}[\boldsymbol{\varepsilon}\mid\boldsymbol{X},\boldsymbol{Z}]=0$ (or even the more relaxed assumption $\mathbb{E}[\boldsymbol{\varepsilon}^{\top}(\boldsymbol{X},\boldsymbol{Z})]=0$) when $\boldsymbol{X},\boldsymbol{Z}$ follow a random design, which allows for heteroskedastic noise and typically appears in the regression coefficient tests with asymptotically valid size control. A common approach to handle exchangeable noise is through the idea of permutation tests (Pitman, 1937a, b, 1938). Recently, Lei and Bickel (2021) implemented this idea to the problem of regression coefficient testing. In their seminal work, the authors proposed a cyclic permutation test that achieved finite-sample size validity under Assumption 1 by exploiting the exchangeability of the noise terms. However, to achieve a size $\alpha$ control, their cyclic permutation test requires that $n/p\geq 1/\alpha-1$. For instance, for a sample size of $n=300$ and a targeting Type-I error rate is $\alpha=0.01$, at most $p=2$ covariates are allowed in $\boldsymbol{X}$. This limits the applicability of their test in large dimensions. In this paper, we consider the more challenging question of finite-sample Type-I error control in setting where $p$ is allowed to be of the same order of magnitude as $n$. We propose a residual permutation test (RPT), a permutation-based approach that performs hypothesis tests by manipulating the empirical residuals after regression adjustment. The proposed test is guaranteed to have the correct Type-I error control whenever $p<n/2$. Moreover, our result is fixed design and does not require any regularity conditions on the design matrix $\boldsymbol{X}$.
@@
In addition to proving its finite-sample size validity, we further analyze the statistical power of the proposed test in the regime where $p$ can be up to a constant fraction of $n$, which we will refer to as the proportional regime in this work, especially when the $\varepsilon_{i}$’s follow a heavy-tailed distribution. It should be noted that just with Assumption 1 we can only guarantee our test to have correct size control, the resulting test may not necessarily have power. Indeed, just with Assumption 1, $b$ may not even be identifiable, so that there is no test that is uniformly valid under the null while still maintainng power against some alternative. For example, in the extreme case where $\boldsymbol{Z}=\boldsymbol{X}\beta^{Z}$ for some $\beta^{Z}\in\mathbb{R}^{p}$, i.e., that $\boldsymbol{Z}$ can be expressed as a linear representation of the column vectors of $\boldsymbol{X}$, then $b$ is not an identifiable parameter anymore, i.e., no test can have any power. As we can also see in (6), under Assumption 1 and with such choice of $\boldsymbol{Z}$, we can always have $\mathbb{P}(\phi\leq\alpha)\leq\alpha$ for all $\alpha\in[0,1]$, no matter how large $b$ is. This means that in order for our test to have power, additional assumption on $\boldsymbol{Z}$ must be imposed. As we will discuss further in Section 2.3, statistical methods with robustness to heavy-tailed data have significant demands in practice (Eklund, Nichols and Knutsson, 2016; Wang, Peng and Li, 2015; Cont, 2001), and has been actively studied in both modern statistics and theoretical computer science communities. Despite its importance, there is a lack of available tools that can handle regression coefficient testing under this proportional regime with heavy-tailed noise. In this paper, we fill this gap by showing that under a suitable modelling assumption of $\boldsymbol{Z}$, when the $\varepsilon_{i}$’s are i.i.d. and have a finite $(1+t)$-th order moment for any $t\in[0,1]$, and that $n/p\geq 3+m$ for some $m>0$, our proposed test is asymptotically powerful whenever the coefficient $b$ is of order at least $n^{-t/(1+t)}$. In proving this result, a crucial step is to establish a concentration bound for projected length of a random vector with independent heavy-tailed components. This concentration bound may be of independent interest for future research on statistical procedures with heavy-tailed noise, and is stated in Corollary 8. We also studied the minimax rate optimality of regression coefficient testing with heavy-tailed noises; and proved that in the presence of heavy-tailed noise with only a finite $(1+t)$-th moment, the $n^{-t/(1+t)}$ order requirement for $b$ is essentially rate-optimal.
@@
Since ANOVA has been used extensively in practical applications, as an independent contribution, we provide a more comprehensive analysis of the ANOVA test. Specifically, while ANOVA can be shown to have finite-sample size validity with spherically symmetric noise, our simulations show that the it can substantially violate the nominal size control under more general noise distributions. At the same time, we propose another permutation-based test: naive residual permutation test (naive RPT), which, like ANOVA, also has valid size control under spherically symmetric noise distribution whenever $p<n$. While naive RPT is still not valid for non-spherically symmetric noises, it does appear to have smaller Type I error violations compared to ANOVA.
@@
  • We propose a new test that has finite-sample size validity with fixed-design linear models and exchangeable noises whenever $p<n/2$.
@@
The rest of this paper is organized as follows. In Section 2, we review existing results in regression coefficient testing, permutation- and randomization-based tests and heavy-tailed data. In Section 3, we provide more studies on the finite-sample properties of ANOVA test with non-Gaussian noises, and propose a new test that is easier to implement and more robust to non-Gaussianity. As ANOVA test has been heavily used in practical applications, we believe this is of independent interest. In Section 4, we present our method, and prove its finite-sample size validity. In Sections 5 and 7, we provide power analysis of RPT and study its minimax rate optimality under some heavy-tailed assumptions. Finally, in Section 8 we provide numerical analysis. In Section 9, we end the manuscript with a discussion.
@@ 2.1 Hypothesis testing of regression coefficients
As emphasized by Lei and Bickel (2021), this is a challenging problem, with a “century long effort” in the statistical community to alleviate the strong assumption of ANOVA. In the context of finite-sample size validity, some representative works include Hartigan (1970); Meinshausen (2015). However, the two methods mentioned above still require the noise to follow certain geometric constraint, which is either symmetric about 0 or rotationally invariant. Lei and Bickel (2021) represented, to the best of our knowledge, the first work that established finite-sample size control with only exchangeable noise. However, as mentioned in the introduction, the cyclic permutation test proposed in Lei and Bickel (2021) requires the dimension of $p$ to be much smaller than $n$ for valid size control, and no corresponding statistical power analysis was provided. Alternative test with less restrictive assumptions on dimension $p$ was proposed in D’Haultfœuille and Tuvaandorj (2024), which relaxes the dimensionality assumption by requiring the rows of $\boldsymbol{X}$ to follow a discrete random distribution with a relatively small number of unique values.
Besides finite-sample size validity, a less demanding criteria for coefficient test is the asymptotic size validity. The idea of permutation or randomization have been heavily used to propose asymptotically valid test; see Section 2.2 for more details. In the high-dimensional regime where $p$ is proportional or even much larger than $n$, debiased / desparsified Lasso was proposed to construct confidence intervals and perform coefficient tests (Zhang and Zhang, 2014; Van de Geer et al., 2014; Javanmard and Montanari, 2014). By invoking 1) certain sparsity conditions on the regression coefficients; 2) some regularity conditions on the design matrix $\boldsymbol{X}$ and 3) sharp tail bounds on the noise variables, debiased / desparsified Lasso is guaranteed to establish asymptotically valid p-value and confidence intervals for regression coefficients. We remark that the additional sparsity assumption on the regression coefficients allow for the dimension $p$ to diverge at a much faster rate than $n$ compared to asymptotic regime studied in the current paper. Other follow up studies include Zhu and Bradic (2018); Bradic et al. (2019); Shah and Bühlmann (2023), to name a few.
@@ 2.2 Permutation- and randomization-based hypothesis tests
Other related applications of permutation tests include sharp null hypothesis tests (Caughey, Dafoe and Miratrix, 2017; Caughey et al., 2023), instrumental variable tests (Imbens and Rosenbaum, 2005), and conditional independence tests (Berrett et al., 2020; Kim et al., 2022).
@@ 2.3 Heavy-tailed data
Since the pioneering work by Catoni (2012), the problem of extracting useful information from heavy-tailed data (or the related adversarially contaminated data) has been an active area of research in mathematical statistics and theoretical computer science literature in the past ten years (Bubeck, Cesa-Bianchi and Lugosi, 2013; Lykouris, Mirrokni and Paes Leme, 2018; Lugosi and Mendelson, 2019; Sun, Zhou and Fan, 2020; Fan, Wang and Zhu, 2021). When we allow the dimension $p$ to grow with $n$, heavy-tailed data has been actively studied in mean estimation  (Lugosi and Mendelson, 2019, 2021), regression coefficient estimation (Wang, 2013; Fan, Li and Wang, 2017; Sun, Zhou and Fan, 2020; Pensia, Jog and Loh, 2024) and covariance matrix analysis (Loh and Tan, 2018; Fan, Wang and Zhu, 2021). The definition of “heavy-tail” may vary across different articles. Among all literature working with heavy-tailed noise, our assumptions are most similar to those in Sun, Zhou and Fan (2020); Bubeck, Cesa-Bianchi and Lugosi (2013), which assume that the noise variables has at most a finite $(1+t)$-th order moments for some $t\in(0,1]$ without any geometric or shape constraints. To our knowledge this is also the weakest heavy tail assumption studied in the literature.
@@
3 Finite-sample size validity of ANOVA beyond Gaussianity
@@
Table 1 summarizes the performance of ANOVA test from $100000$ Monte Carlo simulations. We consider the sizes of the ANOVA test at nominal levels $\alpha=0.01,0.05$. According to the simulation results, when the noises of $\boldsymbol{e}$ and $\boldsymbol{\varepsilon}$ follows a standard normal distribution, the ANOVA test has the correct size control, which is consistent with Lemma 1. However, when normality is violated, the ANOVA test will be overly optimistic, with an empirical size more than twice as large as the nominal level in some $1\%$-level tests (this issue is more pronounced if we consider a $0.5\%$ test level; see Table A2 in the Supplementary Material). In particular, the performance of noise type $t_{1}$ is in general worse than that of $t_{2}$, this means that ANOVA test is more vulnerable to heavy-tailed noises. Moreover, the performance of ANOVA is worse with a heavy-tailed design matrix $\boldsymbol{X}$.
@@
It is worth noting that when $\beta=0$ in (1), we can easily construct a permutation test with valid size control by comparing the correlation of $Y$ to $Z$ and to its permutations. From this intuition, a straightforward approach is to first regress both $\boldsymbol{Y}$ and $\boldsymbol{Z}$ onto $\boldsymbol{X}$ to eliminate the influence of $\boldsymbol{X}$, and then to use regression residuals for permutation test construction. Specifically, let $\hat{\boldsymbol{R}}_{\varepsilon}:=(\boldsymbol{I}-\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top})\boldsymbol{Y}$ and $\hat{\boldsymbol{R}}_{e}:=(\boldsymbol{I}-\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top})\boldsymbol{Z}$ be the regression residuals after projecting $\boldsymbol{Y}$ and $\boldsymbol{Z}$ onto $\boldsymbol{X}$ respectively. Let $\boldsymbol{V}_{0}\in\mathbb{R}^{n\times(n-p)}$ be a matrix with orthonormal columns spanning an $(n-p)$-dimensional subspace of $\mathrm{span}(\boldsymbol{X})^{\perp}$, then $\boldsymbol{I}-\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}=\boldsymbol{V}_{0}\boldsymbol{V}_{0}^{\top}$. Hence under $H_{0}:b=0$, the regression residuals $\hat{\boldsymbol{R}}_{\varepsilon}$ satisfy $\hat{\boldsymbol{R}}_{\varepsilon}=\boldsymbol{V}_{0}\boldsymbol{V}_{0}^{\top}\boldsymbol{Y}=\boldsymbol{V}_{0}\boldsymbol{V}_{0}^{\top}\boldsymbol{\varepsilon}$. From above, we construct a test, which we call as naive residual permutation test, based on the projected residuals $\hat{\boldsymbol{\varepsilon}}:=\boldsymbol{V}_{0}^{\top}\hat{\boldsymbol{R}}_{\varepsilon}=\boldsymbol{V}_{0}^{\top}\boldsymbol{Y}$ and $\hat{\boldsymbol{e}}:=\boldsymbol{V}_{0}^{\top}\hat{\boldsymbol{R}}_{e}=\boldsymbol{V}_{0}^{\top}\boldsymbol{Z}$ as
@@
where the $\boldsymbol{P}_{k}\in\mathbb{R}^{(n-p)\times(n-p)}$’s are random permutation matrices that are sampled uniformly at random from the set of all permutation matrices. Lemma 2 shows that under a slightly weaker condition than Lemma 1, $\phi_{\mathrm{naive}}$ has valid Type-I error control.
@@
While Lemma 2 is slightly less stringent than Lemma 1, it still requires the spherical symmetry in distributions. To better understand their empirical performances, we also show the performance of $\phi_{\mathrm{naive}}$ with non-Gaussian noises or non-Gaussian designs in Table 1 and Figures 1(d)-(f) . Without the strong Gaussianity or spherically symmetry assumption, $\phi_{\mathrm{anova}}$ is also not guaranteed to have finite-sample size validity. Nevertheless, when both tests are invalid, the size of naive permutation test is closer to the correct level than its competitor. This indicates that naive test is more robust to non-Gaussian distributions. Moreover, the naive test is an intuitive method and is easy to implement. Thus, the naive test could be used as a preferrable alternative to ANOVA in real data analysis when $n/2\leq p<n$.